HIGH-RESOLUTION NEAR-INFRARED SPECTROSCOPY OF DEUTERATED CH_{2}^{+}

\underline{HAIMING WANG} , CHRISTOPHER P. MORONG, and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

Recent observations of highly deuterated molecules such as ND_{3} and D_{2}CO in prestellar cores and their explanation as due to the extraordinarily high deuterium fractionation of H_{3}^{+} to H_{2}D^{+}, HD_{2}^{+} and D_{3}^{+},^{a} have revealed the importance of observing other deuterated variants of fundamental molecular ions that also play pivotal roles in interstellar chemistry. We have launched an infrared project to study such ions systematically in order to provide their approximate rotational constants for millimeter wave spectroscopists.

We are presenting our work on the near-infrared spectroscopy of CHD^{+} and CD_{2}^{+}. While our search for interstellar CH_{2}^{+} based on our infrared^{b} and near-infrared^{c} laboratory spectra has not been successful due to its extremely high reactivity, its detection in the future is expected in diffuse clouds since it is the intermediate between the abundant CH^{+} and yet to be observed but very important CH_{3}^{+}. CH_{2}^{+} and its deuterated species are also of special interest for theoretical study because of their unique intramolecular dynamics, i.e., the Renner-Teller interaction and quasi-linearity.

Using He-dominated liquid-N_{2} cooled plasmas (∼10 Torr) containing a small amount (∼0.1 Torr) of CH_{4}, CH_{2}D_{2}, CD_{4} and their mixtures, we are searching for the spectra of CD_{2}^{+} and CHD^{+} in the near-infrared from 10,500 cm^{−1} to 12,500 cm^{−1} with our Ti:sapphire laser spectrometer that combines velocity modulation and phase modulation with heterodyne detection for near shot-noise-limited sensitivity. Our search is based on the ab initio calculation by Bunker, Jensen and colleagues^{d} which predicts the Ã(0, 5, 0)^{1} and Ã(0, 4, 0)^{1} ← X̃(0, 0, 0)^{0} bands of CD_{2}^{+}, and the Ã(0, 4, 0)^{1} ← X̃(0, 0, 0)^{0} band of CHD^{+} as the most intense in the region.

aH. Roberts, E. Herbst, and T.J. Millar, ApJ 591, L41 (2003).

bM. Rösslein, C.M. Gabrys, M.-F. Jagod, and T. Oka, J. Mol. Spectrosc. 153, 738 (1992).

cJ.L. Gottfried and T. Oka, J. Chem. Phys. 121, 11527 (2004).

dG. Osmann, P.R. Bunker, P. Jensen, W.P. Kraemer, Chem. Phys. 225, 33 (1997).